How many different five letter passwords can be formed from the lettersA,B,C,D,E,F,G and H if no religion of letters is allowed?
Answer by stanbon(75887)

How many different five letter passwords can be formed from the letters A,B,C,D,E,F,G and H if no repetition of letters is allowed?
------------
8P5 = 8!/(8-5)! = 8!/3! = 8*7*6*5*4 = 6720 words
--------
Cheers,
Stan H.
---------